# SurvMamba: State Space Model with Multi-grained Multi-modal Interaction for Survival Prediction

### v1

@@ 3.5. Interaction Fusion Mamba
To facilitate cross-modal feature interaction and fusion at different granularities, we introduce an IFM module (Algorithm 2). In IFM, we project features from two modalities and employ gating mechanisms to encourage complementary feature learning from each other while suppressing redundant features. After that, multi-modal features are concentrated to form a fused representation. Cross-modality communications are established at fine-grained and coarse-grained level via cascaded IFM. Through this block, we can obtain fine-grained fused features $H_{f}$ from features ${T^{I}}^{\prime}$ and ${T^{G}}^{\prime}$, and get coarse-grained fused features $H_{c}$ from features $T^{I}$ and $T^{G}$, as follows:
(7) $\begin{split}H_{f}&=\textbf{IFM}([{T^{I}}^{\prime},{T^{G}}^{\prime}])\\ H_{c}&=\textbf{IFM}([T^{I},T^{G}]).\end{split}$
Algorithm 2 The IFM module
1:Input: token sequence $T^{a1}_{l-1}:(B,M,D),T^{a2}_{l-1}:(B,M,D)$
2:Output: token sequence $T_{l}:(B,M,D)$
3:for o in {a1, a2} do
4:     $T^{o}_{l-1}:(B,M,D)\leftarrow\text{Norm}(T^{o}_{l-1})$
5:     $x^{\prime}_{o}:(B,M,E)\leftarrow\text{SiLU}(\text{Conv1d}_{o}(x))$
6:     $B_{o}:(B,M,N)\leftarrow\text{Linear}^{B}_{o}(x^{\prime}_{o})$
7:     $C_{o}:(B,M,N)\leftarrow\text{Linear}^{C}_{o}(x^{\prime}_{o})$
8:     $\Delta_{o}:(B,M,E)\leftarrow\log(1+\exp(\text{Linear}^{\Delta}_{o}(x^{\prime}_{o})+\text{Parameter}^{\Delta}_{o}))$
9:     $\overline{A_{o}}:(B,M,E,N)\leftarrow\Delta_{o}\otimes\text{Parameter}^{A}_{o}$
10:     $\overline{B_{o}}:(B,M,E,N)\leftarrow\Delta_{o}\otimes B_{o}$
11:     $Y_{o}:(B,M,E)\leftarrow\text{SSM}(\overline{A_{o}},\overline{B_{o}},C_{o})(x^{\prime}_{o})$
12:end for
13:$z1:(B,M,E)\leftarrow\text{Linear}^{z}(T^{a1}_{l-1})$
14:$z2:(B,M,E)\leftarrow\text{Linear}^{z}(T^{a2}_{l-1})$
15:$Y^{a1}:(B,M,E)\leftarrow Y^{a1}\odot\text{SiLU}(z2)$
16:$Y^{a2}:(B,M,E)\leftarrow Y^{a2}\odot\text{SiLU}(z1)$
17:$T_{l}:(B,M,D)\leftarrow\text{Linear}^{T}(\text{Cat}(Y^{a1},Y^{a2}))$
18:return $T_{l}$
3.6. Survival Prediction
@@
Recognizing the varying significance of features at different granularities for survival prediction, we employ an adaptive fusion strategy to fuse features across granular levels. Utilizing learned hyper-parameter denoted as $\alpha$, we subsequently derive the final feature set for prognostication with $H_{f}$ and $H_{c}$, as follows:
(8) $H=\alpha H_{f}+(1-\alpha)H_{c}.$
@@
Survival prediction estimates the risk probability of an outcome event before a specific time. For the final multi-modal feature $H^{i}$ of $i$-th patient, we use NLL loss  (Zadeh and Schmid, 2020) as the loss function for optimizing survival prediction, following previous works (Xu and Chen, 2023):
$L_{\text{surv}}\left(\{H^{i},t^{i},c^{i}\}_{i=1}^{N_{D}}\right)=$ $-\sum_{i=1}^{N_{D}}c^{i}\log(f_{\text{surv}}^{i}(t|H^{i}))$
$+(1-c^{i})\log(\{1-f_{\text{surv}}^{i}(t-1|H^{i})\})$
(9) $+(1-c^{i})\log(\{1-f_{\text{hazard}}^{i}(t|H^{i})\})$
where $N_{D}$ represents the number of samples in the training set, $f_{\text{surv}}^{i}(t|H^{i})=P(T=t|T\geq t,H^{i})$ denotes the hazard function characterizing the probability of death, and $f_{\text{surv}}^{i}(t|H^{i})=\prod_{k=1}^{t}(1-f_{\text{hazard}}^{i}(k|H^{i}))$ is defined as the survival function, representing the probability of survival up to the time point $t$.
@@
Table 1. c-index (mean ± std) over five TCGA datasets. G. and H. refer to genomic modality (transcriptomics) and histological modality (WSI), respectively. The best results and the second-best results are highlighted in bold and in underline. Cat refers to concatenation, KP refers to Kronecker product.
@@
4. Experiments
In this section, we first describe the datasets and evaluation metrics employed in our study. The experimental results indicate that our approach surpasses contemporary methods in both performance and computational efficiency. Further, we conduct ablation studies to examine the influence of critical components. Finally, from a statistical perspective, we utilize Kaplan-Meier survival curves and the Logrank test to illustrate the effectiveness of survival analysis.
4.1. Datasets and Settings
Datasets.
To demonstrate the performance of our proposed method, we conducted a series of experiments using five public cancer datasets from The Cancer Genome Atlas (TCGA) [^1] http://www.cancer.gov/tcga, which include paired diagnostic WSIs and transcriptomic data alongside verified survival outcomes. The datasets encompass Breast Invasive Carcinoma (BRCA) with 869 cases, Bladder Urothelial Carcinoma (BLCA) with 359 cases, Colon and Rectum Adenocarcinoma (COADREAD) with 296 cases, Uterine Corpus Endometrial Carcinoma (UCEC) with 480 cases, and Lung Adenocarcinoma (LUAD) with 453 cases. Regarding transcriptomic data, we identify 352 unique genomic functions and 42 biological processes, as cataloged in the Kyoto Encyclopedia of Genes and Genomes database (KEGG) [^2] https://www.genome.jp/kegg/. More details about the dataset are provided in Supplementary Material.
Evaluation Metrics.
The models are evaluated using the concordance index (c-index), where a higher value indicates better performance. This index quantifies the proportion of all possible pairs of observations for which the model accurately predicts the sequence of actual survival outcomes.
Implementation.
For WSI preprocessing, we segmented each WSI into regions of $4096\times 4096$ pixels at a magnification level of 20×20$\times$20 ×, subsequently subdividing these regions into patches of $256\times 256$ pixels. The Rectified Adam (RAdam) optimizer was utilized to facilitate model optimization, with a batch size set to 1, a learning rate of $2\times 10^{-4}$, and a weight decay parameter of $5\times 10^{-3}$. The Pathology Encoder (Wang et al., 2021) generates 768-dimensional embeddings that are projected to 512 dimensions. The Genomics Encoder comprises a two-layer feed-forward network designed to produce genomic tokens with 512 dimensions. All computational experiments were conducted on an NVIDIA-A800 GPU. To enhance the robustness of model training, 5-fold cross-validation was applied on all models.
4.2. Comparisons with the State-of-the-Art
To perform a comprehensive comparison with our method, we implemented and evaluated some latest survival prediction methods. We compared our method against the unimodal baselines and the multi-modal SOTA methods. Table 1 shows the experimental results of all methods on all five TCGA datasets.
Unimodal baselines. For transcriptomic data, we specifically implemented SNN (Klambauer et al., 2017) and SNNTrans (Klambauer et al., 2017; Kleinbaum and Klein, 1996). SNN takes the concatenated genomic profiles as a feature vector and predicts the survival outcomes. SNNTrans initially categorizes the genomic data according to the functions and then utilizes TransMIL (Shao et al., 2021) to predict the overall survival. For histology, we compared the SOTA MIL methods ABMIL (Ilse et al., 2018), CLAM (Lu et al., 2021), TransMIL (Shao et al., 2021).
Multi-modal baselines. We compared SOTA methods for multi-modal survival outcome prediction with the previous set-based network architectures (ABMIL, CLAM, TransMIL) with concatenation (Cat) and Kronecker product (KP), two common late fusion mechanisms to integrate bag-level WSI features and genomic features as multi-modal baselines. We also compared MCAT (Chen et al., 2021b), MOTCat (Xu and Chen, 2023) and SurvPath (Jaume et al., 2023) three SOTA methods for multi-modal survival outcome prediction.
@@ 4.2. Comparisons with the State-of-the-Art
Multi-modal SOTA v.s. SurvMamba. SurvMamba outperforms all multi-modal approaches with an overall c-index performance increase ranging from 1.6% to 10.3%. In comparative analyses within each dataset, SurvMamba achieved the highest c-index performance in four out of five cancer benchmarks, demonstrating its potential as a general method for survival prediction task. When compared with MIL-based multi-modal methods with different fusion methods, SurvMamba achieved a performance increase in overall c-index ranging from 5.9% to 10.3%, highlighting the effectiveness of the proposed multi-modal integration method. Besides, our model also outperforms other SOTA multi-modal learning methods by a obvious margin, including MCAT, MOTCat, and SurvPath, showcasing its outstanding ability for multi-modal learning for prognosis.
Table 2. Computational complexity analysis on a single NVIDIA A800 GPU.
Model Params FLOPs c-index
TransMIL (KP) 2339.20K 6.01G 0.658
MCAT 4869.64K 24.796G 0.670
MOTCat 4869.64K 24.796G 0.686
SurvPath 854.15K 5.19G 0.701
SurvMamba 398.85K 2.51G 0.717
Table 3. We compare the effects of different components on the performance (c-index) of SurvMamba on 5 datasets.
@@
Fine-grained Coarse-grained Fine-grained Coarse-grained
@@
Computational Cost Comparison. Existing frameworks for multi-modal survival prediction are based on the attention mechanism to model the intra- and inter-relationships, while our method leverages the state space model. To validate the computational efficiency of our method, we compared SurvMamba with some multi-modal SOTA methods which show the top 5 on the overall c-index of the five TCGA datasets. The computational complexity is evaluated using model parameters (Params) and floating-point operation count (FLOPs). Params evaluates the network’s scale, while FLOPs assess the model’s complexity. As shown in Table 2, we report the overall c-index on the TCGA dataset, Params, and FLOPs for different SOTA models. Compared to TransMIL, the model size and GPU memory of SurvMamba are reduced by 82.9% and 58.2% with a 5.9% better overall c-index. Compared to the second best method SurvPath, SurvMamba also decreases by 53.3% and 51.6% in model size and GPU memory with a 1.6% better c-index. These results show that our model achieves superior performance compared to the state-of-the-art methods with reduced computational cost.
4.3. Ablation Study
In this ablation study, we conducted some extra experiments to investigate the effectiveness of the key components of SurvMamba. Furthermore, we compared the performance of unidirectional and bidirectional Mamba designs in our method.
Ablation of components.
We conducted ablation studies on five TCGA cancer datasets to validate the effectiveness of the proposed modules. Detailed experimental setups are as follows:
@@ 4.3. Ablation Study
Figure 3. Ablation study on the bidirectional design in our proposed SurvMamba.
(a) BRCA
(b) BLCA
(c) COADREAD
(d) UCEC
(e) LUAD
Figure 4. Kaplan-Meier survival curves of SurvMamba on five TCGA cancer datasets, where all patients are stratified into a low-risk group (yellow) and a high-risk group (blue) according to predicted risk scores. Shared areas within two groups refer to the confidence intervals, and a $p$-value of less than 0.05 indicates significant statistical difference.
Impact of Bidirectional and Unidirectional SSM
Figure 3 demonstrates that integrating bidirectionality into state space models significantly enhances prognostic predictions, as indicated by a higher median c-index and a narrower interquartile range. This effect is particularly pronounced in the BRCA dataset, where the boxplots show minimal overlap, suggesting a robust improvement. This improvement is likely due to the model’s enhanced ability to capture complex, bidirectional dependencies within histological and genomic features.
4.4. Survival Analysis
To further validate the effectiveness of SurvMamba for survival analysis, we divided all patients into a low-risk group and a high-risk group based on the median value of the predicted risk scores generated by SurvMamba. Then, we utilize Kaplan-Meier analysis to visualize the survival events of all patients. Meanwhile, we also employ the Logrank test ($p$-value) to measure the statistical significance between the low-risk group and the high-risk group. A $p$-value of less than 0.05 indicates statistical significance. As shown in Figure 4, patients in the low-risk and high-risk groups are stratified clearly on all datasets, demonstrating the prognostic value of SurvMamba in predicting patient outcomes and guiding treatment decisions.
5. Conclusion

### v2

@@ 3.5 Interaction Fusion Mamba
To facilitate cross-modal feature interaction and fusion at different granularities, we introduce an IFM module, which is shown in the middle-right part of Figure 2. In IFM, we project features from two modalities and employ gating mechanisms to encourage complementary feature learning from each other while suppressing redundant features. After that, multi-modal features are concentrated to form a fused representation. Cross-modality communications are established at fine-grained and coarse-grained level via cascaded IFM. Through this block, we can obtain fine-grained fused features $H_{f}$ from features ${T^{I}}^{\prime}$ and ${T^{G}}^{\prime}$, and get coarse-grained fused features $H_{c}$ from features $T^{I}$ and $T^{G}$, as follows:
$\begin{split}H_{f}&=\textbf{IFM}([{T^{I}}^{\prime},{T^{G}}^{\prime}])\\ H_{c}&=\textbf{IFM}([T^{I},T^{G}]).\end{split}$ (7)
3.6 Survival Prediction
@@
Recognizing the varying significance of features at different granularities for survival prediction, we employ an adaptive fusion strategy with learnable weight denoted as $\alpha$ to fuse features across granular levels. We initialize $\alpha$ at 0.5 and include it in the optimizer parameters, adjusting it during optimization to minimize the loss function. Subsequently, we derive the final feature set for prognostication with $H_{f}$ and $H_{c}$, as follows:
$H=\alpha H_{f}+(1-\alpha)H_{c}.$ (8)
@@
Survival prediction estimates the risk probability of an outcome event before a specific time. For the final multi-modal feature $H^{i}$ of $i$-th patient, we use NLL loss  Zadeh and Schmid (2020) as the loss function for optimizing survival prediction, following Xu and Chen (2023).
4 Experiments
4.1 Datasets and Settings
Datasets. To demonstrate the performance of our proposed method, we conducted a series of experiments using five public cancer datasets from The Cancer Genome Atlas (TCGA) [^1] http://www.cancer.gov/tcga, which include paired diagnostic WSIs and transcriptomic data alongside verified survival outcomes. The datasets encompass Breast Invasive Carcinoma (BRCA) with 869 cases, Bladder Urothelial Carcinoma (BLCA) with 359 cases, Colon and Rectum Adenocarcinoma (COADREAD) with 296 cases, Uterine Corpus Endometrial Carcinoma (UCEC) with 480 cases, and Lung Adenocarcinoma (LUAD) with 453 cases. Regarding transcriptomic data, we identify 352 unique genomic functions and 42 biological processes, as cataloged in the Kyoto Encyclopedia of Genes and Genomes database (KEGG) [^2] https://www.genome.jp/kegg/.
Evaluation Metrics. The models are evaluated using the concordance index (c-index), where a higher value indicates better performance. This index quantifies the proportion of all possible pairs of observations for which the model accurately predicts the sequence of actual survival outcomes.
Implementation. For WSI preprocessing, we segmented each WSI into regions of $4096\times 4096$ pixels at a magnification level of 20×20$\times$20 ×, subsequently subdividing these regions into patches of $256\times 256$ pixels. The RAdam optimizer was utilized to facilitate model optimization, with a batch size set to 1, a learning rate of $2\times 10^{-4}$, and a weight decay parameter of $5\times 10^{-3}$. The Pathology Encoder Wang et al. (2021) generates 768-dimensional embeddings that are projected to 512 dimensions. The Genomics Encoder comprises a two-layer feed-forward network designed to produce genomic tokens with 512 dimensions. We set SSM dimension to 16. To enhance the robustness of model training, 5-fold cross-validation was applied on all models. All computational experiments were conducted on an NVIDIA-A800 GPU.
4.2 Comparisons with the State-of-the-Art
Baselines. To perform a comprehensive comparison with our method, we implemented and evaluated some latest survival prediction methods, including: (1) Unimodal baselines. For transcriptomic data, we specifically implemented SNN Klambauer et al. (2017) and SNNTrans Klambauer et al. (2017); Kleinbaum and Klein (1996). For histology, we compared the SOTA MIL methods ABMIL Ilse et al. (2018), CLAM Lu et al. (2021), TransMIL Shao et al. (2021), R2T-MIL Tang et al. (2024). (2) Multi-modal baselines. We compared SOTA methods for multi-modal survival outcome prediction with the previous set-based network architectures (CLAM, TransMIL) with concatenation (Cat) and Kronecker product (KP), two common late fusion mechanisms to integrate bag-level WSI features and genomic features as multi-modal baselines. We also compared CMTA Zhou and Chen (2023), MCAT Chen et al. (2021b), MOTCat Xu and Chen (2023) and SurvPath Jaume et al. (2023) four SOTA methods for multi-modal survival outcome prediction. Table 1 shows the experimental results of all methods on all five TCGA datasets.
@@ 4.2 Comparisons with the State-of-the-Art
Table 1: c-index (mean ± std) over five TCGA datasets. G. and H. refer to genomic modality (transcriptomics) and histological modality (WSI), respectively. The best results and the second-best results are highlighted in bold and in underline.
@@
Multi-modal SOTA v.s. SurvMamba. SurvMamba outperforms all multi-modal approaches with an overall c-index performance increase ranging from 1.6% to 10.3%. In comparative analyses within each dataset, SurvMamba achieved the highest c-index performance in four out of five cancer benchmarks, demonstrating its potential as a general method for survival prediction task. When compared with MIL-based multi-modal methods with different fusion methods, SurvMamba achieved a performance increase in overall c-index ranging from 5.9% to 10.3%, highlighting the effectiveness of the proposed multi-modal integration method. Besides, our model also outperforms other SOTA multi-modal learning methods by a obvious margin, including CMAT, MCAT, MOTCat, and SurvPath, showcasing its outstanding ability for multi-modal learning for prognosis.
Computational Cost Comparison. Existing frameworks for multi-modal survival prediction are based on the attention mechanism to model the intra- and inter-relationships, while our method leverages the state space model. To validate the computational efficiency of our method, we compared SurvMamba with some multi-modal SOTA methods which show the top 5 on the overall c-index of the five TCGA datasets. The computational complexity is evaluated using model parameters (Params) and floating-point operation count (FLOPs). Params evaluates the network’s scale, while FLOPs assess the model’s complexity. As shown in Figure 3, we report the overall c-index on the TCGA dataset, Params, and FLOPs for different SOTA models. Compared to CMTA, the model size and GPU memory of SurvMamba are reduced by 95.70% and 91.65% with a 3.8% better overall c-index. Compared to the second best method SurvPath, SurvMamba also decreases by 53.3% and 51.6% in model size and GPU memory with a 1.6% better c-index. These results show that our model achieves superior performance compared to the state-of-the-art methods with reduced computational cost.
Figure 3: Computational complexity analysis.
4.3 Ablation Study
Ablation of components. We conducted ablation studies to validate the effectiveness of the proposed modules. Detailed experimental setups are as follows:
Table 2: We compare the effects of different components on the performance (c-index) of SurvMamba.
@@ 4.3 Ablation Study
fine coarse fine coarse
@@
Ablation of Mamba. We conducted an ablation study by replacing Mamba modules with Transformer to demonstrate the effectiveness of using the Mamba model for WSIs and transcriptomic data. We measured both the inference runtime and FLOPs for the Transformer-based and Mamba-based approaches. As shown in Table 3, the Mamba-based method demonstrated greater performance with reduced computational costs.
Table 3: We ablate Mamba modules by replacing them with Transformer models.
Model BRCA BLCA COADREAD UCEC LUAD Time Flops
Transformer-based 0.721$\pm$0.016 0.702$\pm$0.046 0.696$\pm$0.021 0.705$\pm$0.052 0.679$\pm$0.035 0.78s 9.73G
Mamba-based 0.737$\pm$0.014 0.720$\pm$0.027 0.697$\pm$0.018 0.731$\pm$0.012 0.702$\pm$0.020 0.22s 2.51G
Impact of Bidirectional and Unidirectional SSM. Figure 4 demonstrates that integrating bidirectionality into state space models significantly enhances prognostic predictions, as indicated by a higher median c-index and a narrower interquartile range. This effect is particularly pronounced in the BRCA dataset, where the boxplots show minimal overlap, suggesting a robust improvement. This improvement is likely due to the model’s enhanced ability to capture complex, bidirectional dependencies within histological and genomic features.
Figure 4: Ablation study on the bidirectional SSM designed in SurvMamba.
4.4 Survival Analysis
To further validate the effectiveness of SurvMamba for survival analysis, we divided all patients into a low-risk group and a high-risk group based on the median value of the predicted risk scores generated by SurvMamba. Then, we utilize Kaplan-Meier analysis to visualize the survival events of all patients. Meanwhile, we also employ the Logrank test ($p$-value) to measure the statistical significance between the low-risk group and the high-risk group. Shared areas within two groups refer to the confidence intervals, and a $p$-value of less than 0.05 indicates significant statistical difference. As shown in Figure 5, patients in the low-risk (yellow) and high-risk (blue) groups are stratified clearly on all datasets, demonstrating the prognostic value of SurvMamba in predicting patient outcomes and guiding treatment decisions.
Figure 5: Kaplan-Meier survival curves of SurvMamba on five TCGA cancer datasets.
5 Conclusion